2(a – 4) + 4 = 5(9 - a)
Answer by checkley77(12844)

2(a – 4) + 4 = 5(9 - a)
2a-8+4=45-5a
2a+5a=45+8-4
7a=49
a=49/7
a=7 ans.
Proof:
2(7-4)+4=5(9-7)
2*3+4=5*2
6+4=10
10=10